Hi,
My question is about finding a measure in simplest form. I have to use the distance formula. The coordinates are (-5,3) and (3,4. The answer I got for that question was 65. Now I have to simplify the answer and I'm not sure what to do.
Thank You
: Hi,
My question is about finding a measure in simplest form. I have to use the distance formula. The coordinates are (-5,3) and (3,4. The answer I got for that question was 65. Now I have to simplify the answer and I'm not sure what to do.
Thank You

Answer by checkley77(3848)

The x distance=3+5=8
The y distance=4-3=1
using the pythagorean theorem a^2+b^2=c^2 we get the distance (c)
8^2+1^2=c^2
64+1=c^2
c^2=65
c=sqrt65
c=8.06226 is the distance from (-5,3) to (3,4)